OpenStudy (anonymous):

Write two equations in standard form that are equivalent to the given equation -7x+6y=-13 and 10x-6y=-12 Can someone tell me how to to do this??? I really want to learn this

OpenStudy (lucaz):

they are already in standard form, so to get an equivalent equation you can multiply it by a number, like 2[-7x+6y=-13] is -14x+12y=-26 the 2nd equation, 1/2[10x-6y=-12] becomes 5x-3y=-6